Assume the average vehicle selling price in the United States last year was $36,420. The average price five years earlier was $31,208. What was the annual increase in the selling price over this time period? Multiple Choice 3.14 percent 3.01 percent 2.56 percent 2.89 percent 1.67 percent
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

36420=31208*(1+r/100)^5

(36420/31208)^(1/5)=(1+r/100)

(1+r/100)=1.0314

r=1.0314-1

=3.14%(Approx).